Consider a 1.245 g sample of glucose (C6H12O6)

a. calculate moles of glucose in the sample

b. Moles of carbon in 1 mole glucose

c. moles and mass of carbon in the sample

d. glucose molecule in the sample


1
Expert's answer
2021-02-22T05:55:57-0500

(a) Molar mass of glucose = "C_6H_{12}O_6 = 6(12)+12(1)+6(16)= 180 g\/mol"


(b) Moles of C in 1 mole of glucose "C_6H_{12}O_6=6" moles


(c) In 180g of sample = 6 moles of carbon

"\rightarrow" In 1g of sample= "\dfrac{6}{180}" moles of carbon

"\rightarrow So," In 1.245g of sample = "\dfrac{6}{180}\times1.245 = 0.0415" moles

and mass of carbon in sample = 0.0415"\times" 12 = 0.498g


(d) Moles of glucose in sample= "\dfrac{1.245}{180}= 6.916\times 10^{-3}moles"

Since, In 1 moles of sample contains 6.023"\times10^{23}molecules"

So, In "6.916\times10^{-3} moles" or 1.245g of sample of glucose contains "6.91\times10^{-3} \times 6.023\times 10^{23}"

"\rightarrow 41.655\times10^{20}=4.1655\times10^{21}" particles